find the roster form from the following set {x:x-2/x-3 = 3,xER}

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
x – 2 = 3 (x – 3)
 
x – 2 = 3x – 9
 
2x = 7
 
x = 3.5
 
hence
 
x = {3.5}
 
Hope it helps
 
Regards
Arun